# Agent A/B: Automated and Scalable Web A/B Testing with Interactive LLM Agents

### v3

@@ Abstract.
A/B testing is central to UI/UX design, yet our formative study with six industry practitioners revealed that it is slowed by scarce user traffic, long run times, and high operational cost. To address these challenges, we introduce Agent A/B, an end-to-end system that deploys large language model (LLM) agents with structured personas to interact with live webpages and generate scalable and behavioral evidence before launch. In a case study on Amazon.com, Agent A/B simulated a between-subjects A/B test of filter panel designs with 1,000 agents and found that the reduced filter list produced more purchases (414 vs. 403), reproducing directional outcomes also observed in a parallel large-scale human experiment. Results further suggest that agent-based simulations can detect interface-sensitive behavioral differences and surface subgroup patterns while offering faster, lower-risk insights. We position Agent A/B as a complement to human testing, enabling earlier prototyping, pre-deployment validation, and hypothesis-driven UX evaluation.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computingccs: Human-centered computing HCI design and evaluation methods
Figure 1. Architecture of Agent A/B system. The system takes user input of agent and testing specifications to automatically 1) generate a large-scale of LLM agents, 2) prepare the testing conditions (e.g., traffic allocation), 3) perform agent-web interaction with the assigned web environment, 4) monitor and record agent behaviors as well as web interactions, and 5) conduct post-testing analysis so that the results will be returned to the end users.:
@@ 1. Introduction
A/B testing—also known as online controlled experimentation—has become a cornerstone methodology for evaluating design decisions in modern web applications (Kohavi et al., 2009, 2007; Kohavi and Longbotham, 2015; Goswami et al., 2015). By comparing alternative interface designs or feature functionalities with real users randomly assigned to control and treatment groups, A/B testing provides rigorous, data-driven evidence for user experience optimization (Johari et al., 2017; Wang et al., 2019; Duan et al., 2021). Major technology companies,including Microsoft (Machmouchi et al., 2017), Yahoo (Tagami et al., 2014), Netflix (Amatriain, 2013), and Amazon (Hill et al., 2017), rely on hundreds of such tests each year to refine features, improve engagement, and guide strategic design choices. Despite its ubiquity, however, traditional A/B testing is often slowed by practical constraints, including heavy reliance on live user traffic, long turnaround times for statistically reliable results, and significant engineering and organizational overhead (Kohavi et al., 2013; Xu et al., 2015; Fabijan et al., 2018).
@@
To better understand these challenges, we conducted a formative study with six industry practitioners who regularly run A/B tests in their day-to-day workflows. Their accounts highlight three recurring bottlenecks (Fig 2). First, there is a lack of lightweight methods for piloting design hypotheses before committing to large-scale traffic allocation. Second, user traffic itself is a scarce and contested resource, particularly for niche products or teams with competing priorities. Third, feedback cycles are slow: weeks may pass between experiment setup and actionable insights, discouraging exploration of alternative designs. Collectively, these constraints mean that many promising ideas are never tested, and early-stage prototyping rarely benefits from rigorous, behavioral-data grounded evaluation.
@@
Recent advances in large language model (LLM)–based agents offer a promising way to address these limitations (Chen et al., 2025). LLM agents have demonstrated the ability to role-play user personas, generate plausible sequences of decisions, and perform multi-step interactions in domains such as e-commerce (Yao et al., 2022), booking (Zhou et al., 2023), and information search (Koh et al., 2024). Yet most prior work has focused on single-session tasks or constrained sandbox environments (Li et al., 2024; Luo et al., 2023; Park et al., 2023; Hua et al., 2023; Wu et al., 2023; Yao et al., 2023), limiting their relevance for evaluating the dynamic and diverse behaviors seen on real-world websites. What remains missing is a scalable framework that can deploy large numbers of persona-driven agents into live web environments, allowing designers to test interface variants and obtain early behavioral signals before committing real user traffic.
@@
We address this gap by proposing Agent A/B (Fig. 1), an end-to-end system for simulating large-scale online experiments using LLM-based agents. The system is designed to be interoperable, able to plug and play with existing agent stacks (e.g., Claude computer-use agents or ReAct (Yao et al., 2023)). As shown in Fig. 1, Agent A/B integrates five coordinated components, spanning agent specification, testing configuration, live agent–web interaction, behavioral monitoring, and automated post-testing analysis. At the core of this pipeline lies the LLM agent sub-system (Fig. 3), which iteratively parses dynamic webpages into structured representations, reasons about the next action given its persona and intention, and grounds that action into concrete DOM operations on live sites. This design enables thousands of agents to run distributed sessions at scale, producing rich traces and statistical comparisons without manual intervention.
Our position is that LLM agents should not replace real user testing; instead, they complement limited piloting sessions by offering earlier, faster, and lower-risk insights for design optimization.
@@
We demonstrate the system through a case study on Amazon.com, where we simulated a between-subjects A/B test comparing the existing full filter panel design against a reduced filter panel novel design powered by a similarity-based ranking algorithm. Using 1,000 agents (500 per condition), we observed that the reduced filter list led to more purchases (414 vs. 403), a modest but statistically reliable increase. Average spending showed a small upward trend, though not statistically significant, and exploratory subgroup analyses suggested stronger effects for male and older personas. Importantly, these agent-based outcomes aligned directionally with the results of a parallel large-scale human A/B experiment, lending confidence to the system’s ability to surface meaningful behavioral signals.
@@
    Agent A/B, a system that deploys LLM-based agents to conduct scalable, persona-driven A/B testing on live webpages.
@@
    Empirical evidence from a case study on Amazon.com showing alignment between agent simulations and large-scale human A/B results.
@@
    Future design direction, demonstrating how agent-based simulations can extend A/B testing toward earlier prototyping, inclusive piloting, and automated design exploration.
@@ 2. Related Work
We situate our work at the intersection of A/B testing, user behavior modeling, autonomous agents, and HCI research on interactive prototyping and simulation. We review four relevant threads of work: (1) limitations of traditional A/B testing practices, (2) advances in automated and optimized experimentation workflows, (3) LLM-based agent simulation across domains, and (4) interactive agents in web environments.
2.1. Limitations of Traditional A/B Testing in Practice
A/B testing has become a foundational methodology of comparing two versions of interfaces and functionalities of a webpage, app, or other digital assets to enable data-driven decisions about design strategies based on user behavior differences (Kohavi et al., 2009, 2007). The application of A/B testing spans both industry and academia, where organizations (e.g., major companies like Microsoft (Machmouchi et al., 2017), Yahoo (Tagami et al., 2014), Netflix (Amatriain, 2013), and Amazon (Hill et al., 2017)) routinely rely on A/B testing to streamline the design, development, and deployment of products.
However, numerous studies in HCI communities have identified structural limitations in this widely adopted methodology. For instance, Fabijan et al. (2017) conducted an in-depth study across software-intensive organizations and found that A/B testing introduces slow iteration loops, costly feature development, and high failure rates—particularly when hypothesis formulation is weak. These findings were reinforced in a broader empirical survey by Fabijan et al. (2018), who emphasized the lack of actionable early-stage insights during the experimentation process. These works collectively underscore the HCI community’s recognition that while A/B testing provides rigor and promising benefits, it falls short in flexibility, speed, and insight generation.
Figure 2. The workflow of web A/B testing and three challenges reported from formative study: 1) the cost and difficulty of securing considerable user traffic for significant results, 2) the whole A/B testing period can span across weeks and months, and 3) limited testing opportunites. :
2.2. Automated Experimentation and Interface Evaluation Tools
Given the aforementioned limitations, the HCI community has grown significant interest in systems that accelerate interface experimentation (Lasecki et al., 2015; Hartmann et al., 2006; Kuznetsov et al., 2022; Swearngin et al., 2012). For example, Apparition (Lasecki et al., 2015) and d.tools (Hartmann et al., 2006) allowed designers to rapidly prototype for physical interaction such as touch interfaces. Fuse (Kuznetsov et al., 2022) enabled the rapid creation of context-aware UI behaviors from demonstrations.
On the experimentation side, Gilotte et al. (2018) proposed techniques for offline A/B testing, using logged user interaction data to estimate counterfactual outcomes for unobserved experimental variants. This reduces the need for live deployments but depends heavily on rich user logs and pre-existing infrastructure. Tamburrelli and Margara (2014) reframed A/B testing as a search problem within a design space, suggesting that evolutionary algorithms could be used to automate variant generation. These systems aim to reduce human effort in running and prioritizing experiments, yet still rely on historical data or user deployment. In contrast, our work explores agent-driven simulation as a complementary mechanism to automatically evaluate designs without live user data.
2.3. User Behavior Simulation: From Cognitive Models to LLM Agents
A longstanding tradition in HCI focuses on simulating user behavior through cognitive models such as GOMS and ACT-R (Card, 2018; Olson and Olson, 1995; John and Kieras, 1996; Gray and Altmann, 2001). Building upon these traditional approaches, a number of works have leveraged inverse reinforcement learning (IRL) to infer user intentions from observed behaviors. Ziebart et al. (2009) introduce a probabilistic approach to IRL, enabling the modeling of complex user strategies in dynamic environments. However, these models are labor-intensive, require domain expertise, and do not generalize easily to open-ended tasks like online shopping. More recent work uses data-driven models to simulate user behavior in online environments. For example, Paranjape et al. (2016) used server logs to reconstruct navigation paths and optimize hyperlink structures. Oppenlaender et al. (2020) introduced CrowdUI, which used inverse reinforcement learning to infer task strategies from real user traces.
With the rise of large language models (LLMs), researchers have begun to explore how LLMs can simulate human behaviors across complex domains (Ge et al., 2023; Wan et al., 2024; Wu et al., 2024). Park et al. (2023) built a generative society of agents simulating daily social behavior in a simulated town. Horton (2023) evaluated whether LLMs could replicate outcomes from real behavioral experiments on prosociality. Furthermore, Lu et al. (2025a) demonstrates that fine-tuning large language models (LLMs) on real-world behavioral data significantly enhances their ability to generate accurate user actions. These systems suggest that LLMs can produce high-fidelity behavioral simulations across diverse tasks. Our work extends this direction into the visually rich, dynamic domain of live web interfaces, focusing on shopping behavior and automated A/B testing experiments.
2.4. LLM Agents in Web Environments
A growing body of work has investigated how autonomous agents can operate within web-based environments, driven by advancements in both machine learning and HCI communities. For instance, WebShop (Yao et al., 2022) introduced a shopping benchmark with templated webpages for studying goal-directed navigation. WebArena (Zhou et al., 2023) extended this to multi-domain service tasks and became a standard evaluation for web agent development. VisualWebArena (Koh et al., 2024) incorporated vision-language models to parse visual cues in interface design. Similarly, WebVoyager (He et al., 2024) and WILBUR (Lutz et al., 2024) emphasized task generalization and adaptive memory in agents interacting with open-ended web content.
However, all these systems operate within simulated environments, which, while useful for reproducibility, abstract away many complexities of the real webpage, such as dynamic loading, layout noise, unexpected modals, and inconsistent structure. Furthermore, the focus of existing work is primarily on task completion, not on evaluating or comparing the user experience across different design configurations. This gap reveals a growing need for systems that can adapt to the diversity and complexity of real-world web environments in which designers often have to balance trade-offs between design constraints and usability. Addressing this gap requires new system architectures and evaluation frameworks that can support interaction in the wild and enable comparative analysis of design decisions in terms of agent efficiency, robustness, and user-centered outcomes.
Participant Gender Job Title Team
P1 F Project Manager PM Team
P2 M Software Development Manager Engineer Team
P3 F Product Manager PM Team
P4 M Product Manager PM Team
P5 M Product Manager PM Team
P6 M Machine Learning Researcher Science Team
Table 1. Demographics of formative study participants.
@@ 3.1. Formative Study Method
To inform the design of the LLM agent-based A/B testing system, we conducted a formative study with six professionals experienced in designing and running A/B tests in industry settings. Through semi-structured interviews, we asked participants to recount a recent A/B test: how it began and ended, the stages involved, technologies used, and challenges encountered. We also probed how they addressed issues and why those strategies were chosen. These interviews helped us understand current practices, challenges, and user needs, which later informed our design decisions for the system.
We used a snowball sampling strategy (Biernacki and Waldorf, 1981) to recruit participants who currently work in the e-commerce industry and have experience designing and conducting A/B tests. We started with our friends, colleagues, and connections who fit the selection criteria and then asked them to refer their connection to participate in our study. All interviewees were located in the United States; A total of 6 participants (four product managers, one software development manager, and one machine learning researcher) were recruited to participate in the interview study (see Table 1 for details).
Interviews were conducted remotely, with audio recorded upon consent. Sessions lasted 45–62 minutes. We transcribed and de-identified all recordings before conducting a grounded theory analysis (Muller and Kogan, 2012). Two co-authors independently performed open coding on practices, tools, and challenges. Through collaborative discussion, we developed a shared codebook and organized codes into categories (e.g., experiment design, system bottlenecks), which were then synthesized into broader themes such as the A/B testing lifecycle and opportunities for system support. We finally applied the developed set of codes to the whole corpus of transcripts.
@@ 3.2. Formative Study Findings
In this section, we first report the overall workflow of an A/B testing project lifecycle (Fig.2), followed by the challenges and coping strategies reported by industrial practitioners. We conclude this section with design requirements for an LLM-Agent based simulation system for A/B testing.
@@
The A/B testing project life cycle, as reported in the interview data, unfolds across seven interdependent stages (Fig.2). It begins with new feature ideation, where ideas originate from individuals or teams and can range from minor UI changes to novel functionalities. Once proposed, the idea enters a team alignment and buy-in stage, involving collaborative discussion and refinement, culminating in approval from higher-level decision-makers. With alignment secured, the A/B testing experiment design stage follows, where critical aspects such as user segmentation, control/treatment conditions, success metrics, and evaluation criteria are meticulously planned—this stage precedes any development to ensure methodological rigor. Next, the feature deployment and iteration stage involves cross-functional collaboration to build and refine the feature, often demanding the most time and resources. Upon completion, the A/B test experiment launch integrates the feature into a live A/B testing environment where user exposure is randomized and monitored. After the test concludes, post-experiment data analysis is conducted with data scientists to assess whether success criteria were met and to understand the underlying reasons. Finally, the feature decision is made based on the outcome: successful features are launched, while unsuccessful ones are shelved.
Participants reported that the whole process can take anywhere from 3 months to a year. As aforementioned, the A/B test process is definitely a team sport, where different job roles need to collaborate with each other to design and develop the feature, and to design, launch, and analyze the A/B test experiment.
Challenges and Coping Strategies
Interviewees described several challenges in the current A/B testing workflow. One major issue is the high cost of new feature development, which often requires multiple engineers working for months to produce a functional version. During this lengthy process, it’s difficult to gather early user feedback to iterate on the design. The only formal feedback comes through the A/B test experiment, which may be too late. Some feature owners use colleagues as alpha testers, but this feedback is typically biased. Thus, interviewees expressed a need for lightweight prototyping and early-stage user testing methods that can offer feedback before committing to a formal A/B test.
“We had features … required changes across all [UI] stacks and beyond in other services, and it took one year and a half [to develop]…” (P2, Software Development Manager)
Another common challenge is competing for user traffic. When multiple teams want to test features that affect similar UI components, their experiments must be serialized, as running them in parallel would interfere with results. Currently, prioritization happens late in the process after development and just before the test launch, which could significantly delay the following steps. Interviewees called for a more data-driven way to manage and prioritize experiment launches.
“The biggest pain point, I think, is ‘traffic lanes’ … So experiments look at the same area; you cannot run them in parallel, so we have to split the [user] traffic. But we don’t have an automated way to split the traffic properly… ” (P2, Software Development Manager)
Some teams reported that up to one-third of experiments fail to meet their predefined success criteria. This high failure rate is frustrating because A/B testing in its current form leads to a one-shot decision: if a feature misses even one metric, the opportunity is lost. Interviewees expressed interest in predictive tools that could help forecast experiment outcomes so that they can revise features or experimental designs before the launch.
“The hypothesis behind [a failed feature design] that is it’s easier to access [the UI element when put it at the bottom] … and to match our competitors. [A/B test experiment shows] customer hates it, [because] they can no longer find the UI element that they used to and they quit. We saw the abandonment rate go through the roof [in the A/B test experiment].” (P3, Product Manager)
In summary, while A/B testing is widely adopted and supported by infrastructure such as user-routing platforms, interviewees identified several shortcomings: the high development cost, the lack of user traffic prioritization, and frequent experiment failure. They called for innovations that support early-stage prototyping, internal evaluation without real user exposure, and predictive assessment of test outcomes before the actual A/B testing period starts.

### v4

# Agent A/B: Automated and Scalable A/B Testing on Live Websites with Interactive LLM Agents

@@ Abstract.
A/B testing is central to UI/UX design, yet our formative study with six industry practitioners revealed that it is slowed by scarce user traffic, long runtimes, and high operational costs. To address these challenges, we introduce Agent A/B, an end-to-end system that deploys large language model (LLM) agents with structured personas to interact with live webpages and generate scalable behavioral evidence before launch. In a case study on Amazon.com, Agent A/B simulated a between-subjects A/B test of filter panel designs with 1,000 agents and found that the reduced filter list produced more purchases, reproducing directional outcomes also observed in a parallel large-scale human experiment. Results further suggest that agent-based simulations can detect interface-sensitive behavioral differences and surface subgroup patterns while offering faster, lower-risk insights. We position Agent A/B as a complement to human testing, enabling earlier prototyping, pre-deployment validation, and hypothesis-driven UX evaluation.
@@
journalyear: 2026copyright: ccconference: Extended Abstracts of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Extended Abstracts of the 2026 CHI Conference on Human Factors in Computing Systems (CHI EA ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772363.3799039isbn: 979-8-4007-2281-3/2026/04ccs: Human-centered computing Usability testing
Figure 1. Architecture of Agent A/B. From agent and test specifications, the system (1) generates LLM agents, (2, 3) runs controlled web interactions, and (4, 5) analyzes logged behaviors for experiment owners.
A flowchart diagram illustrates the A/B testing workflow enabled by UXAgent. On the left, a system user (e.g., UX researcher or product manager) provides two types of input: LLM Agent specifications (e.g., number of agents, demographic distributions, attributes) and A/B Testing configuration (e.g., feature designs, agent intentions, metrics). These inputs lead to two parallel processes: LLM Agent Generation (Step 1) and Testing Condition Preparation (Step 2). The generated agents are then assigned to either a control or treatment group (Step 3), where they interact with their designated web environments. The control group uses the baseline interface, while the treatment group uses a modified (new) design. Simulation results are collected (Step 4) and passed to the Post-Testing Analysis module (Step 5), which returns A/B Testing Results (e.g., session outcomes, average actions, failure rates) back to the user. The visual includes labeled arrows, icons representing agents and interfaces, and text boxes describing each component.
@@ 1. Introduction
A/B testing (i.e., online controlled experimentation) is a cornerstone for evaluating design decisions in modern web applications (Kohavi et al., 2009, 2007; Kohavi and Longbotham, 2015; Goswami et al., 2015). By comparing alternative interfaces or features with users randomly assigned to control and treatment groups, it provides rigorous, data-driven evidence for user experience optimization (Johari et al., 2017; Wang et al., 2019; Duan et al., 2021). Major technology companies, including Microsoft (Machmouchi et al., 2017), Yahoo (Tagami et al., 2014), Netflix (Amatriain, 2013), and Amazon (Hill et al., 2017), run hundreds of such tests annually to refine features, improve engagement, and guide strategic design choices. Despite its ubiquity, traditional A/B testing is costly, slow to iterate, and difficult to revise when designs are flawed, due to long turnaround times and substantial engineering and organizational overhead (Kohavi et al., 2013; Xu et al., 2015; Fabijan et al., 2018).
@@
Our goal is to better support A/B testing experimenters in designing and running experiments. To ground our system design, we conducted a formative study with six industry practitioners who regularly run large-scale A/B tests. Their accounts reveal three recurring bottlenecks (Fig. 2): (1) limited support for lightweight piloting before large-scale traffic allocation, (2) scarce and contested user traffic, and (3) slow feedback cycles between experiment setup and actionable insights. These constraints leave many promising ideas untested, and early-stage prototyping rarely benefits from rigorous, behaviorally grounded evaluation.
@@
Recent advances in large language model (LLM) agents offer a promising way to address these limitations (Chen et al., 2025b; Zhang et al., 2024; Jia et al., 2024; Zhang et al., 2025b; Chen et al., 2025a; Wang et al., 2025a, b; Zhang et al., 2025a; Wang et al., 2026). LLM agents can role-play user personas and perform multi-step interactions in domains such as e-commerce (Yao et al., 2022), booking (Zhou et al., 2023), and information search (Koh et al., 2024). However, most prior work focuses on single-session tasks or constrained environments (Li et al., 2024; Luo et al., 2023; Park et al., 2023; Hua et al., 2023; Wu et al., 2023; Yao et al., 2023), limiting applicability to real-world websites. What remains missing is a scalable framework that deploys large numbers of persona-driven agents in live web environments to surface early behavioral signals before allocating real user traffic.
@@
We address this gap with Agent A/B (Fig. 1), an end-to-end system for simulating large-scale online experiments with LLM-based agents. The system is interoperable and can plug into existing agent stacks (e.g., Claude computer-use agents or ReAct (Yao et al., 2023)). As shown in Fig. 1, Agent A/B integrates five components: agent specification, testing configuration, live agent–web interaction, behavioral monitoring, and automated analysis. At its core, LLM agents iteratively interpret dynamic webpages and ground persona-driven decisions into concrete DOM actions, enabling thousands of distributed sessions and scalable behavioral comparisons without manual intervention.
We emphasize that LLM-agent-based A/B testing is not a replacement for real user testing, but a complementary tool to help experiment owners (e.g., UX researchers and product managers) mitigate traffic scarcity, slow iteration cycles, and collaboration challenges (Kuang et al., 2022; Feng et al., 2023).
@@
We demonstrate Agent A/B through a case study on Amazon.com, simulating a between-subjects A/B test comparing the existing full filter panel with a reduced, similarity-based design. Using 1,000 agents (500 per condition), the reduced design produced more purchases than the baseline, with a modest but statistically significant increase. Spending showed a small upward trend, and agent outcomes aligned directionally with a parallel large-scale human A/B experiment, suggesting the system can surface meaningful behavioral signals.
@@
    Agent A/B, an end-to-end system for scalable, persona-driven A/B testing with LLM-based web agents.
@@
    Empirical evidence from an Amazon.com case study showing alignment between agent simulations and human A/B results.
@@
    Design implications for using agent-based simulation to support early prototyping and exploratory experimentation.
@@ 2. Related Work
We situate our work at the intersection of A/B testing, user behavior modeling, and HCI research on interactive prototyping and simulation. We review two threads: (1) limitations of traditional A/B testing and advances in automated experimentation workflows, and (2) LLM-based agent simulation across domains.
2.1. Limitations and Automation of A/B Testing
A/B testing is a foundational methodology for comparing interface and feature variants to support data-driven design decisions (Kohavi et al., 2009, 2007), and is widely used in both industry and academia, including by companies such as Microsoft (Machmouchi et al., 2017), Yahoo (Tagami et al., 2014), Netflix (Amatriain, 2013), and Amazon (Hill et al., 2017). Despite its ubiquity, prior HCI research documents persistent limitations, including slow iteration cycles, high development costs, limited early-stage feedback, and methodological challenges such as metric selection, data quality, and statistical reliability under real-world constraints (Fabijan et al., 2017, 2018; Gupta et al., 2019; Larsen et al., 2022). Beyond methodological issues, experimentation is shaped by socio-technical factors: UX practitioners face collaboration overhead, fragmented analyses, time pressure, and limited influence over data and model decisions in AI-intensive products (Kuang et al., 2022; Feng et al., 2023; Zdanowska and Taylor, 2022; Lu et al., 2022). In response, the HCI community has explored tools for rapid prototyping, in-situ sensemaking, and offline or automated experimentation (Lasecki et al., 2015; Hartmann et al., 2006; Kuznetsov et al., 2022; Oppenlaender et al., 2020; Gilotte et al., 2018; Tamburrelli and Margara, 2014), but these approaches still rely on historical data or live user traffic. Our work builds on these efforts by investigating agent-driven simulation as a complementary approach for evaluating interface designs without consuming real user traffic.
2.2. User Behavior Simulation with LLM Agents
HCI has long studied user behavior simulation through cognitive models such as GOMS and ACT-R (Card, 2018; Olson and Olson, 1995; John and Kieras, 1996; Gray and Altmann, 2001), and later through inverse reinforcement learning to infer user intentions from observed behavior (Ziebart et al., 2009; Kangasrääsiö and Kaski, 2018). While these approaches enable fine-grained modeling, they are labor-intensive, require domain expertise, and do not generalize well to open-ended tasks like online shopping; data-driven methods using server logs improve scalability but still rely on historical traces and fixed environments (Paranjape et al., 2016). Recent advances in large language models (LLMs) have renewed interest in behavior simulation across complex domains (Ge et al., 2023; Wan et al., 2024; Wu et al., 2024), with prior work demonstrating LLM agents’ ability to simulate social behavior (Park et al., 2023), reproduce behavioral experiments (Horton, 2023), Agent-as-a-judge evaluation (Chen et al., 2025c), and improve action fidelity when trained on real-world data (Lu et al., 2025). In parallel, web-agent benchmarks such as WebShop (Yao et al., 2022), WebArena (Zhou et al., 2023), VisualWebArena (Koh et al., 2024), WebVoyager (He et al., 2024), and WILBUR (Lutz et al., 2024) have advanced autonomous web interaction, but remain largely confined to simulated or templated environments and focus on task completion rather than comparative user experience evaluation. Our work bridges these lines by deploying persona-driven LLM agents directly on live web interfaces to enable scalable user behavior simulation for A/B testing and early-stage design evaluation.
@@ 3.1. Formative Study Method
To inform the design of our LLM agent-based A/B testing system, we conducted a formative study with six industry professionals experienced in designing and running large-scale A/B tests. We used semi-structured interviews to elicit recent A/B testing experiences, focusing on workflows, tools, challenges, and how participants addressed them. Participants were recruited via snowball sampling (Biernacki and Waldorf, 1981) from the e-commerce industry until thematic saturation. All six participants were based in the United States (four product managers, one software development manager, and one machine learning researcher) and routinely managed A/B tests with over one million users. Interviews were conducted remotely, lasted 45–62 minutes, and were audio-recorded with consent. Recordings were transcribed, de-identified, and analyzed using grounded theory (Muller and Kogan, 2012). Two authors independently coded the data, reconciled differences, and synthesized themes related to A/B testing workflows, bottlenecks, and opportunities for system support.
@@ 3.2. Formative Study Findings
Interview data revealed a seven-stage A/B testing lifecycle (Fig. 2). The process begins with feature ideation, where teams propose new UI changes or functionalities, followed by team alignment and buy-in involving discussion, refinement, and managerial approval. Teams then conduct experiment design, specifying user segmentation, control and treatment conditions, and success metrics prior to development. The feature development and iteration stage follows, requiring cross-functional collaboration and substantial engineering effort. Once implemented, the experiment launch deploys the feature into a live A/B testing environment with randomized user exposure. Afterward, post-experiment analysis interprets outcomes and behavioral drivers, leading to a final feature decision to launch or discard the feature. Participants reported that this end-to-end process typically spans three months to over a year and requires close coordination across multiple roles.
Challenges and Coping Strategies.
Interviewees highlighted recurring challenges in current A/B testing workflows, including the high cost and long timelines of feature development, which leave little opportunity for early user feedback before formal testing. To cope, some teams rely on internal alpha testing with colleagues, though participants noted this feedback is often biased and insufficient for guiding iteration, underscoring the need for lightweight prototyping and early-stage evaluation. Interviewees also described intense competition for user traffic, which forces experiments affecting similar interface components to be serialized and prioritized late in the process, delaying deployment and iteration. Finally, high experiment failure rates often result in one-shot decisions with limited opportunities for refinement, leading participants to call for tools that enable internal evaluation, data-driven prioritization, and predictive assessment of experiment outcomes prior to live A/B testing.